# Yirui Zhou, Xiaowei Liu, Xiaofeng Zhang, Yangchun Zhang

### v1

@@ Abstract
This paper tackles the efficiency and stability issues in learning from observations (LfO). We commence by investigating how reward functions and policies generalize in LfO. Subsequently, the built-in reinforcement learning (RL) approach in generative adversarial imitation from observation (GAIfO) is replaced with distributional soft actor-critic (DSAC). This change results in a novel algorithm called Mimicking Observations through Distributional Update Learning with adequate Exploration (MODULE), which combines soft actor-critic’s superior efficiency with distributional RL’s robust stability.
@@ 1 Introduction
Imitation learning (IL) (Pomerleau,, 1991; Ng et al.,, 2000; Syed and Schapire,, 2007; Ho and Ermon,, 2016), a realm distinct from standard reinforcement learning (RL) (Puterman,, 2014; Sutton and Barto,, 2018), is independent on rewards provided by the environment. This characteristic makes IL particularly suited for numerous real-world applications (Bhattacharyya et al.,, 2018; Shi et al.,, 2019; Jabri,, 2021). The general IL paradigm leverages the guidance from expert demonstrations with information of both states and actions to mimic an outstanding policy (Abbeel and Ng,, 2004; Ho and Ermon,, 2016; Kostrikov et al.,, 2020). According to the strategy of policy training, IL is divided into two main schemes based on policy training strategy: on-policy and off-policy training. The on-policy scheme (Ho and Ermon,, 2016; Chen et al.,, 2020) is noted for its stability but requires a significant volume of samples. Conversely, the traditional off-policy scheme (Kostrikov et al.,, 2019; Zhou et al.,, 2022) is more sample-efficient but is at risk of higher training variance.
@@
Figure 1: An example exhibits the misjudgment by the learned reward. $s$ and $s^{\prime}$ denote the expert state and its next state, respectively. Assume $a_{1}$ be the expert action in the scenario of LfD, where $(s,a_{1})$ induces the same optimal policy as $(s,s^{\prime})$ in LfO. However, the agent may execute another action $a_{i}$ ($i=2,\ldots,n$) but still output $s^{\prime}$ acting on the transition dynamics. This results in the misjudgment.
We find that the direct application of the two primary LfD policy training schemes in LfO tends to magnify their adverse effects.
@@
  • Instability issue in the traditional off-policy scheme worsens with LfO. There exists a lack of training compatibility between the learned reward (depend on $s,s^{\prime}$) and the agent (depend on $s$, action $a$) in LfO. For instance (Fig. 1), the sequences $(s,a_{i},s^{\prime}),i=1,\ldots,n$ generated by the agent will be uniformly recognized as approaching expert. This increases the risk of misjudgment by the learned reward, leading to training instability.
@@
$\hskip-8.5359pt\mathop{\min}_{\pi\in\Pi}\mathop{\max}_{r\in\mathcal{R}}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi}}\left[r(s,s^{\prime})\right],$ (1)
where $\Pi$ and $\mathcal{R}$ represent the policy class and the reward function class, respectively. $\pi_{\rm E}$ denotes the expert policy and $\mu_{\pi}$ is the state transition distribution induced by the policy $\pi$ (Yang et al., 2019a, ; Zhu et al.,, 2020). Consequently, we establish that the generalization results of Eq. (1) align with the findings for generative adversarial imitation learning (GAIL) in (Xu et al.,, 2020, Lemma 2) and (Zhou et al.,, 2023, Theorem 1). This alignment allows us to adapt GAIL’s computational strategy for addressing the LfO computation. Particularly, we modify the policy optimization method in generative adversarial imitation from observation (GAIfO) (Torabi et al., 2018b, ) with distributional soft actor-critic (DSAC) (Ma et al.,, 2020; Duan et al.,, 2022), referred to as the Mimicking Observations through Distributional Update Learning with adequate Exploration (MODULE) algorithm. The benefit of MODULE is two-fold:
@@ 2.1 Generalization properties for learning from demonstrations
A body of research has delved into the generalization capability within the context of LfD (Chen et al.,, 2020; Xu et al.,, 2020; Zhou et al.,, 2023, 2024). Specifically, Chen et al., (2020) showed that it is practicable to ensure the generalization of $\mathcal{R}$-distance by properly controlling the reward function in GAIL with parameterized rewards. In the follow-up study, under the metric of the neural network distance (Arora et al.,, 2017), the connection between generalization capabilities and expected return is established by analyzing the linear span of the discriminator class (Xu et al.,, 2020). Zhou et al., (2023) contributed insights into the policy class, leveraging the state-action distribution error to analyze the generalization of GAIL. Furthermore, (Zhou et al.,, 2024) investigates the behavior of the generalization error bound within the reward transfer paradigm in transfer imitation learning (TIL), where state-action demonstration data is available exclusively in the source environment. In contrast to prior research, this paper explores LfO generalization with respect to the reward function class and the policy class, analyzing each respectively within a single environment.
@@ 2.2 Learning from observations
One effective strategy for tackling LfO is to estimate proxy rewards from the states encountered by both the learning agent and the expert, and then integrate these into a RL framework to guide the agent toward improving its performance by maximizing the total accumulated reward (Torabi et al., 2018b, ; Yang et al., 2019a, ; Lee et al.,, 2021; Jaegle et al.,, 2021; Liu et al.,, 2024). The majority of methods that can derive proxy rewards include those employing the adversarial training mechanism (Torabi et al., 2018b, ; Yang et al., 2019a, ; Sikchi et al.,, 2023), optimal transport (OT) (Villani et al.,, 2009) based approaches (Liu et al.,, 2024; Chang et al.,, 2024), and techniques focused on estimating goal proximity (Lee et al.,, 2021; Bruce et al.,, 2023).
@@
OT based approaches. Recently, the concept of OT (Villani et al.,, 2009) has been incorporated into the field of IL to label the proxy rewards (Liu et al.,, 2024; Chang et al.,, 2024). Liu et al., (2024) designed the automatic discount scheduling (ADS) mechanism to dynamically adjust the discount factor in RL, focusing on earlier rewards and progressively emphasizing later rewards. Observational off-policy Sinkhorn (OOPS) (Chang et al.,, 2024), an LfO method that optimizes a reward function by reducing the discrepancy between the state trajectories of experts and learners, quantified through the Wasserstein distance.
Goal proximity estimation. Lee et al., (2021) utilized the information of task progress readily and intuitively available in demonstration data, i.e., goal proximity, to derive a generalizable reward. In addition, (Bruce et al.,, 2023) introduces the explore like experts (ELE) method, which employs expert data to approximate the monotonically increasing progress function through the temporal difference between frames that serves as an exploration reward, aiding the agent in efficiently navigating sparse-reward environments.
Among these methods, our MODULE algorithm hinges on adversarial training as its foundational methodology. This process develops a function capable of distinguishing the data between the expert and the agent effectively, thereby constructing the proxy reward.
In a distinct area of LfO research, a number of approaches concentrate on model-based techniques, where the agent learns an inverse dynamics model (IDM) to infer the missing action information for the expert (Torabi et al., 2018a, ; Zhu et al.,, 2020; Cândido Ramos et al.,, 2024). Behavioral cloning from observation (BCO) (Torabi et al., 2018a, ) enables the immediate acquisition of imitation policies by leveraging an IDM learned from state trajectories alone, without requiring access to demonstrator actions or post-demonstration interaction with the environment. Off-policy learning from observations (OPOLO) (Zhu et al.,, 2020) uses principled off-policy improvement and accelerates the learning process by employing an inverse action model to guide policy updates, thereby promoting distribution matching in the mode-covering perspective. Mimicking better by matching the approximate action distribution (MAAD) (Cândido Ramos et al.,, 2024) incorporates an IDM within the framework of on-policy LfO. Our MODULE algorithm stands distinct from model-based approaches, yet it has the potential to augment stability when incorporating the distributional update strategy with model-based components.
@@ 3 Preliminaries
The model of how the agent interacts with the environment is captured using a Markov decision process (MDP) characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r_{gt},\gamma)$ with the finite state space $\mathcal{S}$ and the action space $\mathcal{A}$. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the transition dynamics, $r_{gt}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the ground truth reward, and $\gamma$ represents the discount factor. A stochastic policy $\pi(a|s)$ assigns probabilities to actions $a\in\mathcal{A}$ based on the current state $s\in\mathcal{S}$. Following previous works (Xu et al.,, 2020; Yang et al., 2019a, ; Zhu et al.,, 2020), we use three notations for the stationary distributions associated with $\pi$: $d_{\pi}(s)$ for the state, $\rho_{\pi}(s,a)$ for the state-action pair and $\mu_{\pi}(s,s^{\prime})$ for the state transition. Their definitions are listed in Table 1.
@@
[…]
@@ 3.1 Adversarial training in learning from observations
In comparison to LfD, the challenge in LfO increases due to the fact that only states, rather than state-action pairs, are provided as expert guidance. The adversarial training problem in LfO is formulated as the min-max optimization problem in Eq. (1). The objective is to find a policy within the policy class $\Pi$ that recovers the expert policy $\pi_{\rm E}$ using expert observation data. The learned reward is designed to distinguish between state transition pairs from expert observations and those produced by the agent. Simultaneously, the policy aims to generate interaction data that challenges the learned reward’s ability to make this distinction, which corresponds to solving the following problem
$\mathop{\max}_{\pi\in\Pi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi}}\left[r(s,s^{\prime})\right]$
under a fixed reward $r\in\mathcal{R}$.
@@ 3.2 Distributional soft actor-critic
[…]
@@ 4.1 Generalization for the reward function
We denote $d_{\mathcal{R}}^{\rm LfO}(\hat{\mu}_{\pi},\hat{\mu}_{\pi^{\prime}})$ as the empirical LfO reward distance, where $\hat{\mu}_{\pi}$ and $\hat{\mu}_{\pi^{\prime}}$ are the empirical counterparts of $\mu_{\pi}$ and $\mu_{\pi^{\prime}}$, respectively, estimated from $n$ samples. In practice, Eq. (1) achieves the minimization of the empirical LfO reward distance between $\hat{\mu}_{\pi_{\rm E}}$ and $\hat{\mu}_{\pi}$. We now introduce a theorem that elucidates the generalization capability for the reward function of LfO, which is similar to the result for GAIL in (Xu et al.,, 2020, Lemma 2).
@@ 4.2 Generalization for the policy
In practice, the shape of the learned reward $r$ is diverse, with possible value ranges including (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), $[0,+\infty)$, $(-\infty,0]$, $[0,1]$, etc (Wang and Li,, 2021). Here we use a non-negative $r$ as a representative example to investigate the generalization property for the policy in LfO.
@@
where $\hat{c}_{r}^{(m)}=\sum_{i=1}^{m}r(s^{(i)},s^{\prime(i)})$ is the empirical state transition reward coefficient and ^$\hat{\mathbb{R}}(s^{(i)},s^{\prime(i)})=r(s^{(i)},s^{\prime(i)})/\hat{c}_{r}^{(m)}$ denotes the empirical counterpart of $\mathbb{R}$ with $m$ samples. Next, we provide the generalization capability for the policy of LfO, which is similar to the consequence of GAIL in (Zhou et al.,, 2023, Theorem 1).
@@ 5 MODULE algorithm
$\hskip-19.91684pt\mathop{\min}_{\theta}\mathop{\max}_{\phi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r_{\phi}(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\theta}}}\left[r_{\phi}(s,s^{\prime})\right],$ (3)
@@
$\mathop{\min}_{\theta}\mathop{\max}_{\phi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r_{\phi}(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\theta}}}\left[r_{\phi}(s,s^{\prime})\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\alpha\mathbb{H}(\pi_{\theta})-\frac{\mu}{2}\|\phi\|_{2}^{2},$ (4)
@@
which is essentially the objective of maximum entropy RL. On this basis, our MODULE algorithm is designed to combine SAC (adjusting entropy automatically) with the technique of distributional RL during policy updates. Particularly, (1) SAC, grounded in the principle of maximum entropy RL, guarantees that the learned policy accounts for all beneficial actions (Haarnoja et al., 2018b, ). (2) Distributional RL, known for its powerful ability to model the distribution over returns and capture randomness, improves computational stability (Bellemare et al.,, 2017).
@@
    $J_{Q}(w)=\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}^{\rm I}}\Big{[}\frac{1}{2}\big{(}Q_{w}^{\rm soft}(s_{t},a_{t})-(r_{\phi}(s_{t},s_{t+1})+\gamma\mathbb{E}_{s_{t+1}\sim p(\cdot|s_{t},a_{t})}[V_{\bar{w}}^{\rm soft}(s_{t+1})])\big{)}^{2}\Big{]},$
@@
    $\hskip 2.84544ptV_{\bar{w}}^{\rm soft}(s_{t})=\mathbb{E}_{a_{t}\sim\pi_{\theta}}[Q_{\bar{w}}^{\rm soft}(s_{t},a_{t})-\alpha\log(\pi_{\theta}(a_{t}|s_{t}))].$
@@
  • The objective of the policy network parameter $\theta$ is to minimize the expected KL divergence in Eq. (2) (Haarnoja et al., 2018a, ; Haarnoja et al., 2018b, ), i.e., minimizing
    $\hskip-25.6073ptJ_{\pi}(\theta)=\mathbb{E}_{s_{t}\sim\mathcal{D}^{\rm I}}\big{[}\mathbb{E}_{a_{t}\sim\pi_{\theta}}[\alpha\log\pi_{\theta}(a_{t}|s_{t})-Q_{w}^{\rm soft}(s_{t},a_{t})]\big{]}.$ (6)
@@
    $\hskip-8.5359ptJ(\alpha)=\mathbb{E}_{a_{t}\sim\pi_{\theta}}\big{[}-\alpha\log\pi_{\theta}(a_{t}|s_{t})-\alpha H_{0}\big{]},$ (7)
@@
    $\hskip-2.84544pt\mathcal{L}_{\kappa}\left(\delta_{ij}\right)=\begin{cases}~{}~{}~{}~{}~{}~{}~{}~{}~{}\frac{1}{2}\delta_{ij}^{2},&\text{if}\left|\delta_{ij}\right|\leq\kappa,\\ \kappa\left(\left|\delta_{ij}\right|-\frac{1}{2}\kappa\right),&\text{otherwise},\end{cases}$
@@
    $Q_{w}^{\rm soft}(s,a)=\min_{i=1,2}Q_{w_{i}}^{\rm soft}(s,a),Q_{w_{i}}^{\rm soft}(s,a)=\Psi\left[Z_{\tau,w_{i}}(s,a)\right],$ (9)

### v2

# On Generalization and Distributional Update for Mimicking Observations with Adequate Exploration

@@ Abstract
Learning from observations (LfO) replicates expert behavior without needing access to the expert’s actions, making it more practical than learning from demonstrations (LfD) in many real-world scenarios. However, directly applying the on-policy training scheme in LfO worsens the sample inefficiency problem, while employing the traditional off-policy training scheme in LfO magnifies the instability issue. This paper seeks to develop an efficient and stable solution for the LfO problem. Specifically, we begin by exploring the generalization capabilities of both the reward function and policy in LfO, which provides a theoretical foundation for computation. Building on this, we modify the policy optimization method in generative adversarial imitation from observation (GAIfO) with distributional soft actor-critic (DSAC), and propose the Mimicking Observations through Distributional Update Learning with adequate Exploration (MODULE) algorithm to solve the LfO problem. MODULE incorporates the advantages of (1) high sample efficiency and training robustness enhancement in soft actor-critic (SAC), and (2) training stability in distributional reinforcement learning (RL). Extensive experiments in MuJoCo environments showcase the superior performance of MODULE over current LfO methods.
@@ 1 Introduction
Imitation learning (IL) (Pomerleau,, 1991; Ng et al.,, 2000; Syed and Schapire,, 2007; Ho and Ermon,, 2016), a realm distinct from standard reinforcement learning (RL) (Puterman,, 2014; Sutton and Barto,, 2018), does not rely on rewards provided by the environment. This characteristic makes IL well-suited for various real-world applications (Bhattacharyya et al.,, 2018; Shi et al.,, 2019; Jabri,, 2021). The general IL paradigm leverages the guidance from expert demonstrations with information of both states and actions to mimic an outstanding policy (Abbeel and Ng,, 2004; Ho and Ermon,, 2016; Kostrikov et al.,, 2020). According to the policy training strategy, IL is divided into two main schemes based on policy training strategy: on-policy and off-policy training. The on-policy scheme (Ho and Ermon,, 2016; Chen et al.,, 2020) is noted for its stability but requires a significant volume of samples. Conversely, the traditional off-policy scheme (Kostrikov et al.,, 2019; Zhou et al.,, 2022) is more sample-efficient but is at risk of higher training variance.
@@
Figure 1: An illustration of the misjudgment by the learned reward: Given $(s,s^{\prime})$ as the expert’s state $s$ and its next state $s^{\prime}$ in LfO, which corresponds to the expert state-action pair $(s,a_{1})$ in LfD, misjudgment occurs when the agent might execute alternative actions $a_{i}$ ($i=2,\ldots,n$) that have the same transition dynamics leading to $s^{\prime}$.
We find that directly applying the two primary LfD policy training schemes in LfO tends to magnify their adverse effects.
@@
  • Instability issue in the traditional off-policy scheme worsens with LfO. There exists a lack of training compatibility between the learned reward (depending on states $s,s^{\prime}$) and the agent (depending on state $s$, action $a$) in LfO. For instance (Fig. 1), the sequences $(s,a_{i},s^{\prime}),i=1,\ldots,n$ generated by the agent will be uniformly recognized as approaching the expert. This increases the risk of misjudgment by the learned reward, leading to training instability.
@@
$\thinspace\mathop{\min}_{\pi\in\Pi}\mathop{\max}_{r\in\mathcal{R}}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi}}\left[r(s,s^{\prime})\right],$ (1)
where $\Pi$ and $\mathcal{R}$ represent the policy class and the reward function class, respectively. $\pi_{\rm E}$ denotes the expert policy and $\mu_{\pi}$ is the state transition distribution induced by the policy $\pi$ (Yang et al., 2019a, ; Zhu et al.,, 2020). Consequently, we establish the generalization results of Eq. (1), which are based on the findings of generative adversarial imitation learning (GAIL) in (Xu et al.,, 2020, Lemma 2) and (Zhou et al.,, 2023, Theorem 1). This enables us to handle the computation involved in LfO using GAIL’s computational strategy. In particular, we modify the policy optimization method in generative adversarial imitation from observation (GAIfO) (Torabi et al., 2018b, ) with distributional soft actor-critic (DSAC) (Ma et al.,, 2020; Duan et al.,, 2022), referred to as the Mimicking Observations through Distributional Update Learning with adequate Exploration (MODULE) algorithm. The benefit of MODULE is two-fold:
@@ 2.1 Generalization properties for learning from demonstrations
A body of research has delved into the generalization capability within the context of LfD (Chen et al.,, 2020; Xu et al.,, 2020; Zhou et al.,, 2023, 2024). Specifically, Chen et al., (2020) showed that it is practicable to ensure the generalization of $\mathcal{R}$-distance by properly controlling the reward function in GAIL with parameterized rewards. In the follow-up study, under the metric of the neural network distance (Arora et al.,, 2017), the connection between generalization capabilities and expected return is established by analyzing the linear span of the discriminator class (Xu et al.,, 2020). Zhou et al., (2023) contributed insights into the policy class, leveraging the state-action distribution error to analyze the generalization of GAIL. Furthermore, (Zhou et al.,, 2024) investigates the behavior of the generalization error bound within the reward transfer paradigm in transfer imitation learning (TIL), where state-action demonstration data is available exclusively in the source environment. Nevertheless, LfO generalization with respect to both the reward function class and the policy class within a single environmental setting remains an under-explored area in the literature.
@@ 2.2 Learning from observations
One effective strategy for tackling LfO is to estimate proxy rewards from the states encountered by both the learning agent and the expert, and then integrate these into a RL framework to guide the agent toward improving its performance by maximizing the total accumulated reward (Torabi et al., 2018b, ; Yang et al., 2019a, ; Lee et al.,, 2021; Jaegle et al.,, 2021; Liu et al.,, 2024). The majority of methods that can derive proxy rewards include optimal transport (OT) (Villani et al.,, 2009) based approaches (Liu et al.,, 2024; Chang et al.,, 2024), techniques focused on estimating goal proximity (Lee et al.,, 2021; Bruce et al.,, 2023), and those employing the adversarial training mechanism (Torabi et al., 2018b, ; Yang et al., 2019a, ; Sikchi et al.,, 2023).
OT based approaches. Recently, the concept of OT (Villani et al.,, 2009) has been incorporated into the field of IL to label the proxy rewards (Liu et al.,, 2024; Chang et al.,, 2024). Liu et al., (2024) designed the automatic discount scheduling (ADS) mechanism to dynamically adjust the discount factor in RL, focusing on earlier rewards and progressively emphasizing later rewards. Observational off-policy Sinkhorn (OOPS) (Chang et al.,, 2024) optimizes a reward function by reducing the discrepancy between the state trajectories of experts and learners, quantified through the Wasserstein distance.
Goal proximity estimation. Lee et al., (2021) utilized the information of task progress readily and intuitively available in demonstration data, i.e., goal proximity, to derive a generalizable reward. In addition, (Bruce et al.,, 2023) introduces the explore like experts method, which employs expert data to approximate the monotonically increasing progress function through the temporal difference between frames that serves as an exploration reward, aiding the agent in efficiently navigating sparse-reward environments.
@@
Among these methods, our MODULE algorithm hinges on adversarial training as its foundational methodology. This process develops a function capable of effectively distinguishing the data between the expert and the agent, thereby constructing the proxy reward.
In a distinct area of LfO research, several approaches concentrate on model-based techniques, where the agent learns an inverse dynamics model (IDM) to infer the missing action information for the expert (Torabi et al., 2018a, ; Zhu et al.,, 2020; Cândido Ramos et al.,, 2024). Indeed, behavioral cloning from observation (BCO) (Torabi et al., 2018a, ) can imitate policies without requiring access to demonstrator actions or post-demonstration interaction with the environment. However, BCO demonstrates poor performance with a learned IDM and fails to learn the consistent policy across both state-based and image-based experiments (Sikchi et al.,, 2024). An alternative approach named off-policy learning from observations (OPOLO) (Zhu et al.,, 2020) uses principled off-policy improvement and accelerates the learning process by employing an inverse action model to guide policy updates, thereby promoting distribution matching in the mode-covering perspective. Mimicking better by matching the approximate action distribution (Cândido Ramos et al.,, 2024) incorporates an IDM within the framework of on-policy LfO. Our MODULE algorithm differs from model-based approaches, yet it has the potential to augment stability when incorporating the distributional update strategy with model-based components.
@@ 3 Preliminaries
The model of how the agent interacts with the environment is captured using a Markov decision process (MDP) characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r_{gt},\gamma)$ with the finite state space $\mathcal{S}$ and the action space $\mathcal{A}$. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the transition dynamics, $r_{gt}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the ground truth reward, and $\gamma$ represents the discount factor. A stochastic policy $\pi(a|s)$ assigns probabilities to actions $a\in\mathcal{A}$ based on the current state $s\in\mathcal{S}$. Following previous works (Xu et al.,, 2020; Yang et al., 2019a, ; Zhu et al.,, 2020), we use the notations for the stationary distributions associated with $\pi$: $d_{\pi}(s)$ for the state, $\rho_{\pi}(s,a)$ for the state-action pair and $\mu_{\pi}(s,s^{\prime})$ for the state transition. Their definitions are listed in Table 1.
@@
State transition distribution $\mu_{\pi}(s,s^{\prime})$ $\mathcal{S}\times\mathcal{S}$ $\int_{\mathcal{A}}\rho_{\pi}(s,a)\mathcal{P}(s^{\prime}|s,a)da$
@@ 3.1 Adversarial training in learning from observations
Due to the fact that only states rather than state-action pairs are provided as expert guidance, the challenge in LfO increases in comparison to LfD. The adversarial training problem in LfO is formulated as the min-max optimization problem in Eq. (1). It aims to find a policy within the policy class $\Pi$ that recovers the expert policy $\pi_{\rm E}$ using expert observation data. The learned reward is designed to distinguish between state transition pairs from expert observations and those produced by the agent. Simultaneously, the policy aims to generate interaction data that challenges the learned reward’s ability to make this distinction, which corresponds to solving the following problem under a fixed reward $r\in\mathcal{R}$:
$\mathop{\max}_{\pi\in\Pi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi}}\left[r(s,s^{\prime})\right].$
@@ 3.2 Distributional soft actor-critic
$\mathcal{T}_{DS}^{\pi}Z(s,a):\overset{D}{=}r(s,a)+\gamma\big[Z(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\big],s^{\prime}\sim\mathcal{P}(\cdot|s,a),a^{\prime}\sim\pi(\cdot|s^{\prime}),$
where $U:\overset{D}{=}V$ specifies identical probability distributions for random variables $U$ and $V$. Then, the distributional soft policy improvement stage updates the policy towards the exponential form of the soft action-value function:
@@ 4.1 Generalization for the reward function
We denote $d_{\mathcal{R}}^{\rm LfO}(\hat{\mu}_{\pi},\hat{\mu}_{\pi^{\prime}})$ as the empirical LfO reward distance, where $\hat{\mu}_{\pi}$ and $\hat{\mu}_{\pi^{\prime}}$ are the empirical counterparts of $\mu_{\pi}$ and $\mu_{\pi^{\prime}}$, respectively, estimated from $n$ samples. In practice, Eq. (1) achieves the minimization of the empirical LfO reward distance between $\hat{\mu}_{\pi_{\rm E}}$ and $\hat{\mu}_{\pi}$. We now introduce a theorem that elucidates the generalization capability for the reward function of LfO, which is based on the result for GAIL in (Xu et al.,, 2020, Lemma 2).
@@ 4.2 Generalization for the policy
In practice, the shape of the learned reward $r$ is diverse, with possible value ranges including $(-\infty,+\infty)$, $0,+\infty)$, $(-\infty,0]$, $[0,1]$, etc (Wang and Li,, 2021). Here, we use a non-negative $r$ as a representative example to investigate the generalization property for the policy in LfO.
@@
where $\hat{c}_{r}^{(m)}=\sum_{i=1}^{m}r(s^{(i)},s^{\prime(i)})$ is the empirical state transition reward coefficient and $\hat{\mathbb{R}}(s^{(i)},s^{\prime(i)})=r(s^{(i)},s^{\prime(i)})/\hat{c}_{r}^{(m)}$ denotes the empirical counterpart of $\mathbb{R}$ with $m$ samples. Next, we provide the generalization capability for the policy of LfO, which is based on the consequence of GAIL in (Zhou et al.,, 2023, Theorem 1).
@@ 5 MODULE algorithm
$\thinspace\mathop{\min}_{\theta}\mathop{\max}_{\phi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r_{\phi}(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\theta}}}\left[r_{\phi}(s,s^{\prime})\right],$ (3)
@@
$\mathop{\min}_{\theta}\mathop{\max}_{\phi}\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}\left[r_{\phi}(s,s^{\prime})\right]-\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\theta}}}\left[r_{\phi}(s,s^{\prime})\right]-\alpha\mathbb{H}(\pi_{\theta})-\frac{\mu}{2}\|\phi\|_{2}^{2},$ (4)
@@
which is essentially the objective of maximum entropy RL. Driven by this, our MODULE algorithm is designed to combine SAC (which adjusts entropy automatically) with the technique of distributional RL during policy updates. Particularly, (1) SAC, grounded in the principle of maximum entropy RL, guarantees that the learned policy accounts for all beneficial actions (Haarnoja et al., 2018b, ). (2) Distributional RL, known for its powerful ability to model the distribution over returns and capture randomness, improves computational stability (Bellemare et al.,, 2017).
@@
    $J_{Q}(w)=\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}^{\rm I}}\Big[\frac{1}{2}\big(Q_{w}^{\rm soft}(s_{t},a_{t})-(r_{\phi}(s_{t},s_{t+1})+\gamma\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}[V_{\bar{w}}^{\rm soft}(s_{t+1})])\big)^{2}\Big],$
@@
    $V_{\bar{w}}^{\rm soft}(s_{t})=\mathbb{E}_{a_{t}\sim\pi_{\theta}}[Q_{\bar{w}}^{\rm soft}(s_{t},a_{t})-\alpha\log(\pi_{\theta}(a_{t}|s_{t}))].$
@@
  • The objective of the policy network parameter $\theta$ is to minimize the expected Kullback-Leibler (KL) divergence in Eq. (2) (Haarnoja et al., 2018a, ; Haarnoja et al., 2018b, ), i.e., minimizing
    $\thinspace J_{\pi}(\theta)=\mathbb{E}_{s_{t}\sim\mathcal{D}^{\rm I}}\big[\mathbb{E}_{a_{t}\sim\pi_{\theta}}[\alpha\log\pi_{\theta}(a_{t}|s_{t})-Q_{w}^{\rm soft}(s_{t},a_{t})]\big].$ (6)
@@
    $\thinspace J(\alpha)=\mathbb{E}_{a_{t}\sim\pi_{\theta}}\big[-\alpha\log\pi_{\theta}(a_{t}|s_{t})-\alpha H_{0}\big],$ (7)
@@
    $\thinspace\mathcal{L}_{\kappa}\left(\delta_{ij}\right)=\begin{cases}~~~~~~~~~\frac{1}{2}\delta_{ij}^{2},&\text{if}\left|\delta_{ij}\right|\leq\kappa,\\ \kappa\left(\left|\delta_{ij}\right|-\frac{1}{2}\kappa\right),&\text{otherwise},\end{cases}$
@@
    $Q_{w}^{\rm soft}(s,a)$ $=\min_{i=1,2}Q_{w_{i}}^{\rm soft}(s,a),$
    $Q_{w_{i}}^{\rm soft}(s,a)$ $=\Psi\left[Z_{\tau,w_{i}}(s,a)\right],$ (9)